OpenStudy (anonymous):

in the following relation, identify the domain and range. {(2.-4), (-2,4), (2,5), (5,-1), (-1,5)}

OpenStudy (ash2326):

Domain is the first element of each set Range is the second element of each set Can you try now?

OpenStudy (anonymous):

I understand how to get it, but i can't find the answer.

OpenStudy (ash2326):

Write the first element of each set {2, -2, 2, 5, -1} remove the repeated elements {2, -2, 5, -1} Write in ascending order {-2, -1, 2, 5} This is the domain

OpenStudy (anonymous):

thanks.

OpenStudy (ash2326):

Can you do the same thing to find range? here take the second element

OpenStudy (anonymous):

i Still don't understand

OpenStudy (ash2326):

Which part you have doubt?

OpenStudy (anonymous):

Range.

OpenStudy (anonymous):

I had the domain.

OpenStudy (ash2326):

Please Read my posts again :)

OpenStudy (anonymous):

How would the graphs of f(x) = 2x + 6 and g(x) = 2x + 3 compare if graphed on the same coordinate plane?

OpenStudy (anonymous):

ignore that i put the wrong one. Given f(x) = 3x + 2 and g(x) = -2x - 4, find h(x) = f(x) - g(x).

OpenStudy (ash2326):

\[h(x)=f(x)-g(x)\] plugin the values of f(x) and g(x) \[h(x)=(3x+2)-(-2x-4)\] Can you simplify this ?